OpenStudy (anonymous):

Find the perimeter and area of a right triangle if the shortest side is 12 cm and the longest side is 20 cm. Include correct units with each part of your solution.

OpenStudy (lgbasallote):

area is just \[\frac {bh}2\] base is 12 and height is 20

OpenStudy (anonymous):

Do i multiply?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

240?

OpenStudy (lgbasallote):

now divide by 2

OpenStudy (anonymous):

120

OpenStudy (lgbasallote):

that's your area now

OpenStudy (lgbasallote):

to solve for perimeter... \[P = b + h + \sqrt{b^2 + h^2}\]

OpenStudy (anonymous):

Help?

OpenStudy (lgbasallote):

again... b is 12 and h is 20

OpenStudy (lgbasallote):

just substitute

OpenStudy (anonymous):

P=b+h+b^2+h^2 12+16+20=48 cm. Sides are 12,16,20

OpenStudy (lgbasallote):

wait...im checking....

OpenStudy (lgbasallote):

how did you get 16?